michael bought 3 pencils and 2 rulers at a total cost of $2.30. nacey bought 3 rulers and 2 pencils for $2.00. what would be the cost of one pencil?
Answer by stanbon(75887)

michael bought 3 pencils and 2 rulers at a total cost of $2.30. nancy bought 3 rulers and 2 pencils for $2.00. what would be the cost of one pencil?
--------------------
3p + 2r = 230
2p + 3r = 200
-----------------
9p + 6r = 3*230
4p + 6r = 2*200
--------------------
5p = 290
p = 58 cents (cost of one pencil)
------------
Cheers,
Stan H.
-------------